### v1

# Hydrogen-Alpha as a Tracer of Star Formation in the SPHINX Cosmological Simulations

Hydrogen-Alpha as a Tracer of Star Formation in the SPHINX Cosmological Simulations
The Hydrogen-alpha () emission line in galaxies is a powerful tracer of their recent star formation activity. With the advent of JWST, we are now able to routinely observe in galaxies at high redshifts () and thus measure their star-formation rates (SFRs). However, using “classical” SFR() calibrations to derive the SFRs leads to biased results because high-redshift galaxies are commonly characterized by low metallicities and bursty star-formation histories, affecting the conversion factor between the luminosity () and the SFR. In this work, we develop a set of new SFR() calibrations that allow us to predict the SFRs of -emitters at with minimal error. We use the SPHINX cosmological simulations to select a sample of star-forming galaxies representative of the -emitter population observed with JWST. We then derive linear corrections to the classical SFR() calibrations, taking into account variations in the physical properties (e.g., stellar metallicities) among individual galaxies. We obtain two new SFR() calibrations that, compared to the classical calibrations, reduce the root mean squared error (RMSE) in the predicted SFRs by dex and dex, respectively. Using the recent JWST NIRCam/grism observations of -emitters at , we show that the new calibrations affect the high-redshift galaxy population statistics: (i) the estimated cosmic star-formation density decreases by %, and (ii) the observed slope of the star-formation main sequence increases by .
Key Words.:
galaxies: fundamental parameters – galaxies: evolution – galaxies: formation – galaxies: high-redshift – galaxies: star formation1 Introduction
One of the most fundamental and basic properties of galaxies is the rate at which new stars form, the star formation rate (SFR). The SFR is sensitive to the interplay between the gas inflow and outflow rates of galaxies, and the efficiency at which this gas is converted into stars (e.g., Bouché et al., 2010; Lilly et al., 2013; Behroozi et al., 2019). The SFR can fluctuate on different timescales (e.g., Sparre et al., 2015; Iyer et al., 2020; Flores Velázquez et al., 2021), which range from dynamical timescales of molecular clouds (s of Myrs), through galaxy merger timescales ( Myr), to long (Gyr-scale) coherences in accretion rates (e.g., Matthee & Schaye, 2019; Wang & Lilly, 2020; Tacchella et al., 2020; Wan et al., 2025). Observationally, SFRs can be measured using various diagnostics (e.g., Kennicutt & Evans, 2012; De Looze et al., 2014; Davies et al., 2019) that are either based on the light from massive, short-lived stars (e.g., the rest-frame UV), or processed radiation (e.g., nebular lines in HII regions or dust emission). Each of these diagnostics probes different timescales of star formation and is affected by dust attenuation differently.
At high redshifts (), the rest-frame UV emission used to be the most sensitive SFR indicator that could be measured for large galaxy samples, tracing star formation on the Myr timescales (e.g., Bouwens et al., 2015). However, dust corrections in the rest-frame UV are significant, with potentially up to 50% of the star formation rate density at being obscured by dust (e.g., Zavala et al., 2021; Sun et al., 2025). In contrast, the Hydrogen-alpha () emission – an alternative SFR indicator – is emitted in the rest-frame optical ( Å) and thus has a much lower sensitivity to dust attenuation. The luminosity of the line is directly proportional to the luminosity of ionizing photons that are predominantly emitted by very young stellar populations, making sensitive to star formation on shorter timescales ( Myr; e.g., Leitherer et al., 1999). As the JWST can now measure the distribution of luminosities out to based on spectroscopy and narrow-band imaging (e.g., Clarke et al., 2024; Pirie et al., 2024; Covelo-Paz et al., 2025; Fu et al., 2025), it is timely to reassess the role of the H emission as a tracer of the SFR.
The conversion factor between the H luminosity111Throughout this work, we assume a negligible contribution of active galactic nuclei to the H emission. and the SFR primarily depends on assumptions on the properties of stellar populations, in particular the star formation history (SFH), the initial mass function (IMF), binarity and stellar metallicity. The SFH and the IMF are relevant because the ionizing photon emissivity from a single stellar population drops sharply after a few Myr when massive, short-lived O stars leave the main sequence. Binary interactions can extend these ionizing lifetimes (Götberg et al., 2017), smoothing the otherwise steep drop-off after a few Myrs. At fixed stellar mass and age, the metallicity impacts the effective temperature of stellar atmospheres due to line blanketing effects, affecting the ionizing luminosity as well.
The classical SFR() calibration proposed by Kennicutt (1998) (see also Kennicutt & Evans 2012) assumes a constant SFH over Myr and solar metallicity, based on the typical conditions in galaxies in the low-redshift Universe. However, various spectroscopic observations of galaxies at high redshifts suggest that their SFHs are bursty (e.g., Asada et al., 2024; Endsley et al., 2024; Cole et al., 2025; Carvajal-Bohorquez et al., 2025). Given that the gas-phase oxygen abundances are typically measured to be % solar (Curti et al., 2024; Scholte et al., 2025), and early galaxies are expected to be -enhanced due to the rapid formation timescales (e.g. Chruślińska et al., 2024), stellar metallicities are likely to be very low (), implying hotter stars and a more efficient photon production at a fixed SFR (e.g., Theios et al., 2019), compared to low redshift.
The aim of this paper is to revisit the classical SFR() calibrations to predict the SFRs of galaxies at high redshifts (). Rather than using idealistic models with constant SFHs and fixed stellar metallicities, we use simulated galaxies and their modeled luminosities from the SPHINX cosmological radiation-hydrodynamical simulation (Rosdahl et al., 2018; Katz et al., 2023). These cosmological simulations self-consistently model SFHs and metal enrichment processes, and are calibrated to match various observables at high redshifts (). The simulated volume is well matched to the typical volumes of deep JWST surveys, which enables a high resolution that can capture SFR variations on short timescales ( Myr) and approximate the distribution of gas in the interstellar medium relevant for radiative transfer effects.
In Sect. 2, we briefly describe the SPHINX simulations and explain how the H luminosities are derived. In Sect. 3, we present a set of simulation-inferred SFR() calibrations that predict the SFRs of SPHINX galaxies with minimal error. Finally, we discuss the implications of our work in Sect. 4, and give a summary in Sect. 5.
2 Methods
In this work, we use the simulated galaxy data from SPHINX, a suite of cosmological radiation-hydrodynamic simulations (Rosdahl et al., 2018). SPHINX leverages the adaptive mesh refinement code RAMSES-RT (Rosdahl et al., 2013; Teyssier, 2002) to simulate tens of thousands of galaxies in the epoch of reionization. The ISM is resolved down to co-moving pc, corresponding to the physical resolution of pc at . The dark matter halos are resolved down to the atomic cooling threshold ( M⊙). Hydrodynamics, non-equilibrium thermochemistry, and radiative transfer of Lyman continuum radiation (LyC; Å) are performed self-consistently on-the-fly, with subgrid implementations for gas cooling, star formation, metal enrichment, and feedback. The on-the-fly radiative emission and propagation of LyC photons make SPHINX an excellent model to accurately predict how emission traces star formation in the simulated galaxies.
Our analysis is based on the data from the SPHINX20 Public Data Release, Version 1 (SPDRv1; Katz et al., 2023). SPHINX20 is the largest simulation of the SPHINX suite, with a periodic box with volume of ( cMpc)3 simulated down to a redshift of (Rosdahl et al., 2022). SPHINX20 uses the BPASSv stellar population synthesis (SPS) model (Stanway & Eldridge, 2018) that adopts a Kroupa (2001a) IMF with an upper mass cutoff of 100 M⊙ and a slope of from to M⊙ and from to M⊙. The intrinsic Balmer line luminosities are generated by interpolating the CLOUDY v17.03 models (Ferland et al., 2017) for gas cells that host stellar particles with unresolved Stromgren spheres, or by using the non-equilibrium ionizing fractions from the simulation otherwise. The equivalent widths are calculated by taking into account both the stellar and nebular continuum emission, generated as detailed in Katz et al. (2023). Post-processing of line and continuum radiation, subject to absorption and scattering by dust, is carried out using the radiative transfer code RASCAS (Michel-Dansac et al., 2020) assuming the Small Magellanic Cloud (SMC) dust model from Michel-Dansac et al. (2020) (for details, see Katz et al., 2023).
SPDRv1 only includes SPHINX20 galaxies with M⊙ yr-1, thus excluding objects that are likely too faint to be observed with JWST (Katz et al., 2023). However, an SFR cut leads to incompleteness in the SFR— relation at low luminosities which can bias our results. Therefore, we limit our selection of galaxies from SPDRv1 to those that have an observed luminosity () greater than erg s-1. This limit is motivated by the recent LF measurements from JWST NIRCam/grism that reach similar luminosities at the faint end (e.g., Fu et al., 2025). Crucially, this selection yields a complete -emitter sample, which we check by running RASCAS on all SPHINX20 haloes whose intrinsic luminosity () exceeds the luminosity cut. We also try erg s-1 and erg s-1 cuts, finding little impact on our results. With this additional luminosity cut, our final sample comprises , 66, 48, 36, 22, 17, and 13 galaxies at , 5, 6, 7, 8, 9, and 10, respectively, with a total of galaxies.
The SN feedback model in SPHINX is calibrated to match the UV luminosity function (UVLF) at (see App. C in Rosdahl et al., 2018). A possible caveat is that the stellar mass-halo mass relation in SPHINX is – dex higher than predicted by, e.g., the Behroozi et al. (2019) model that matches a broad array of observational data including stellar mass functions at – and median UV—stellar mass relations at – (see Fig. 4 in Katz et al., 2023). Furthermore, the stellar mass function (SMF) in SPHINX is significantly overestimated compared to the SMF inferred from the JWST/NIRCam observations at – (Weibel et al., 2024). Hence, if stellar masses of SPHINX galaxies are overestimated, the amount of dust in the simulation should be higher than in the real Universe in order to match the UVLF. The overproduction of dust would then explain the fact that SPHINX predicts lower and luminosities at fixed stellar mass than those measured by JWST (e.g., Covelo-Paz et al., 2025; Meyer et al., 2024). Although the possibility of SPHINX galaxies having too much dust does not affect our main results as they are based on the intrinsic properties (Sect. 3), it will become important later for understanding the errors in the SFR() calibrations that use the observed (Sect. 4).
3 Results
In this section, we investigate the intrinsic (i.e., without taking into account dust attenuation) SFR— relation in the SPHINX20 cosmological simulation. We compare this relation to the SFR() calibrations from the literature and study how the variations in the physical properties of individual galaxies (metallicity, stellar age, etc.) impact the accuracy of the predicted SFRs. Based on this analysis, we develop a set of new SFR() calibrations that minimize the error in -derived SFRs at high redshifts.
Unless otherwise specified, we use the SFRs averaged over the last Myr of the star-formation history. This choice is motivated by the fact that nebular emission lines such as trace stars with lifetimes of –10 Myr (e.g., Kennicutt & Evans, 2012). This is further supported by the simulation itself, where the correlation coefficient between the SFR and peaks between Myr (–; see App. A).
3.1 The SFR— relation in SPHINX
Figure 1 shows the intrinsic SFR— relation for the galaxy sample selected from the SPHINX20 simulation as described in Sect. 2. In the same figure, we show different SFR() calibrations proposed in the literature. These include the original K98 calibration (), the K98 calibration converted to the Kroupa IMF (Kroupa, 2001b; ), and two calibrations from Theios et al. (2019, hereafter T19) calculated for a Kroupa-type IMF and solar metallicity (), and the same IMF but subsolar metallicity (; ), respectively. Here, is the conversion factor between and SFR, defined as follows:
Figure 1 reveals that a constant cannot provide an accurate estimate of the SFRs of SPHINX galaxies across the full range. Specifically, the T19 calibration that assumes (the dashed blue line in Fig. 1) matches well the median SFR— relation in SPHINX at high luminosities (). In contrast, the same SFR() calibration at low luminosities () overestimates the SFRs by – dex. Decreasing would result in a better match at low luminosities; however, the SFRs at high luminosities would be underestimated in that case.
We note that the median SFR— relation in SPHINX at low luminosities lies between the two T19 calibrations that differ only in the assumption on metallicity ( vs. ; indicated by the solid and dashed blue lines in Fig. 1, respectively). Motivated by this result, we test whether the variations in metallicity could be a key factor driving the evolution of the median SFR— relation in SPHINX. We split our sample into low- and high- groups ( and , respectively), and calculate the median stellar metallicity222Hereafter, stellar metallicities (and ages) of SPHINX galaxies are weighted by the LyC luminosity. in each group. We obtain and in the low- and high- groups, respectively. In other words, brighter galaxies are, on average, more metal-rich, which reflects the mass-metallicity relation (MZR) in SPHINX (for discussion, see Katz et al., 2023), and likely causes an upturn of the median SFR— relation at high luminosities (see also Fig. 1 where markers are color-coded by metallicity).
Figure 1 further demonstrates that the original K98 calibration (indicated by the dotted orange line) overestimates the SFRs by – dex compared to SPHINX. One of the main reasons for this discrepancy is the difference in the IMFs: K98 uses the Salpeter (1955) IMF, whereas SPHINX uses a Kroupa-type IMF. After rescaling to the Kroupa IMF, the K98 calibration still predicts – dex higher SFRs than the simulation (the dashed orange line in Fig. 1). This difference cannot be eliminated even when using the T19 calibration which, unlike the K98 calibration, assumes the same stellar population synthesis model as SPHINX (i.e., BPASSv; the dashed blue line in Fig. 1). Therefore, the only way to fully reconcile the K98 calibration and the simulation data is to assume a sub-solar metallicity, for some galaxies as low as (the solid blue line in Fig. 1). As discussed in the introduction, such sub-solar metallicity is well in line with the current estimates of the gas-phase metallicity, as well as with the direct estimates of stellar metallicities of galaxies at (Steidel et al., 2016; Cullen et al., 2019), although we caution that those have been derived assuming constant star formation histories which might impact the results (Matthee et al., 2022).
Figure 1 also reveals that the SFR— relation in SPHINX exhibits significant scatter around the median relation (hereafter ). In particular, we find that decreases from dex at to dex at , and has a median dex. This scatter is neglected by the existing SFR() calibrations, leading to errors in the predicted SFRs of individual galaxies. To quantify this effect, we introduce the SFR bias factor, defined as the ratio between the SFR predicted from and the “true” SFR measured in the simulation: . Here, we use the T19 calibration that assumes to calculate SFR() as this calibration provides the best match to the SPHINX data (see Fig. 1). We note that using a different calibration would only change by a constant factor (in logarithmic scale) without affecting any trends observed between and the physical properties of SPHINX galaxies.
We find that decreases with stellar metallicity (; Fig. 2, top panel), in agreement with our earlier analysis of the median SFR— relation in SPHINX. However, we also report a significant scatter in at fixed : dex. This indicates that metallicity is not the only parameter responsible for the variations in the SFRs of SPHINX galaxies with respect to the SFRs predicted from . In particular, we find that shows a negative correlation with stellar age (; Fig. 2, bottom panel), with an average scatter of dex. High (– dex) that is found in galaxies with young stellar populations ( Myr) can be explained by the fact that the existing SFR() calibrations assume constant star formation over a long timescale (typically Myr or Gyr; e.g., Hao et al., 2011). This assumption breaks down in the simulation, where galaxies exhibit very bursty and stochastic star formation histories (Katz et al., 2023).
We also examine the relationship between and stellar mass and find a negative correlation (), likely driven by the positive scaling of stellar mass with metallicity (see Katz et al., 2023 for the discussion of the MZR in SPHINX). Finally, we check the dependence of on the escape fraction of ionizing radiation () and find a nearly flat trend (), indicating no significant correlation. This result is expected since the strength of nebular emission lines is inversely proportional to the escape fraction of ionizing photons, , and SPHINX galaxies have, on average, low (%). This is in agreement with high-redshift observations, where is also estimated to be low ( %; e.g. Mascia et al., 2024).
3.2 Simulation-informed SFR() calibrations
Using the insights gained from our analysis of the SFR— relation in SPHINX, we discuss possible improvements to the existing SFR() calibrations that take the form of Eq. 1. For example, stellar metallicity could be a useful parameter to include in the calibration to reduce the error in the predicted SFR (see Fig. 2). However, measuring stellar metallicity is difficult in practice, therefore we only use quantities that can be measured robustly from galaxy spectra. Specifically, we run an ordinary least squares (OLS) model similar to Eq. 1 but allowing the coefficient to vary. We use the randomly selected of the sample to train the model and the remaining to validate its performance. The resulting SFR() calibration can be written as follows:
where the last term is the effective correction to Eq. 1. We find that the root mean squared error (RMSE) calculated on the training and test datasets is identical ( dex), indicating that the model performs well on unseen data. Compared to the T19 calibration () which yields dex calculated on the full dataset, the RMSE is reduced by dex (see also Fig. 3).
The SFR() calibration given in Eq. 2 corrects the SFRs by a factor that depends on only, which means that is neglected even when using the updated calibration. To further improve on this, we add a correction to Eq. 2 that depends on the equivalent width (). This choice is motivated by the fact that the traces stellar metallicity and age that both drive (Fig. 2). In addition, the is only weakly sensitive to dust attenuation, which implies that dust corrections have a smaller impact on the results compared to when other observables are used (e.g., ). We therefore run an OLS model that includes , using the same training sample as in Eq. 2. The resulting SFR() calibration can be written as follows:
where the last two terms are the effective corrections to Eq. 1. The new calibration yields dex on both the training and test datasets, dex lower than the RMSE calculated using Eq. 2, or dex lower than the RMSE calculated using the T19 calibration (). Figure 3 compares these calibrations as a function of . In particular, this plot shows that the median in the bins is within dex across the entire range of luminosities () when using Eq. 2 or Eq. 3. This is an improvement over the T19 calibration, which overpredicts the SFRs by dex in galaxies with low-to-medium luminosities (). The standard deviation of in the bins (hereafter ) is nearly identical between the T19 calibration and the SFR() calibration given in Eq. 2 ( dex), but it decreases by dex ( dex) when using Eq. 3, especially at medium luminosities ().
4 Discussion
4.1 Implications
The line has long been known as a reliable indicator of the SFR as it traces the ionizing photon luminosity from massive (), short-lived ( Myr) stars (e.g., Kennicutt, 1998). Over the past decades, has been widely employed in numerous spectroscopic (e.g., Shim et al., 2009; Gunawardhana et al., 2013; Nagaraj et al., 2023) and narrow-band imaging (e.g., Ly et al., 2007; Hayes et al., 2010; Sobral et al., 2013) surveys to measure the cosmic star formation history (SFH) up to . With the recent advent of sensitive near-IR spectroscopy at – m, these measurements have been extended to –, in particular thanks to the JWST/NIRCam slitless spectroscopic observations of large samples of -emitters at the same redshifts (e.g., Covelo-Paz et al., 2025; Fu et al., 2025; Lin et al., 2025).
As observations become more accessible at high redshifts, it is crucial to revisit the classical SFR() calibrations (e.g., Kennicutt, 1998) commonly used at . Based on a realistic galaxy simulation, we have shown that such calibrations bias the predicted SFRs under the physical conditions that are characteristic of high-redshift galaxies observed with JWST (i.e., low metallicity and bursty star formation), potentially also affecting population statistics. To quantify this effect, we assess the impact of the metallicity- and age-dependent SFR bias (see Fig. 2) on the cosmic star formation rate density () measurements. Specifically, we convert the dust-corrected luminosity function at presented by Fu et al. (2025) to the star-formation rate function (SFRF) using the T19 calibration () and our new SFR() calibration given in Eq. 2 which, by design, takes into account the evolution of the SFR— relation with metallicity. Following previous studies, we integrate both SFRFs down to yr-1 which corresponds to (Bouwens et al., 2015). We find that the resulting decreases from to , or by %, when using the new calibration. This suggests that the influence of the evolving SFR— relation on is moderate, likely because galaxies where the SFR bias is highest contribute little to . For example, galaxies with contribute less than % to the total . If we integrate the SFRF only up to , then the change in would be more pronounced: would decrease by %.
The SFR bias also impacts the correlations between the SFR and other galaxy properties, such as stellar mass. In Fig. 4, we show the SFR— relation, also known as the galaxy main sequence (Noeske et al., 2007), in the JWST All the Little Things (ALT) survey (Naidu et al. 2024, Di Cesare et al. in prep.). We find that the slope of the SFMS increases by (%) when we use Eq. 3 instead of the T19 calibration to calculate the SFRs. Equation 2 yields similar results, although with a smaller significance (). In addition, we find that the scatter in the SFMS decreases by dex (%) when using Eq. 3, most noticeably at the low-mass end ( dex at ). In contrast, Eq. 2 produces almost the same scatter as the T19 calibration ( dex). We note that the SMFS in Fig. 4 is impacted by selection effects: some low-mass galaxies are not detected in observations due to their faintness, resulting in a shallow SMFS slope. However, proper modeling of these selection biases is out of scope of this paper and will be discussed elsewhere (see Di Cesare et al. in prep.).
4.2 Dust attenuation
In Sect. 3, we ignored the effects of dust attenuation, assuming that the H line measurements (i.e., and ) can be corrected for such effects without errors. This assumption is also usually made for the SFR() calibrations presented in the literature. However, dust attenuation is known to be an important source of systematic error in the -derived SFRs (e.g., Hopkins et al., 2001). This effect likely persists even at as the amount of dust in -emitters at these redshifts is found to be non-negligible (e.g., Covelo-Paz et al., 2025 reported the median mag for a sample of -emitters at ).
Here, we test the possibility of deriving SFRs directly from the observed emission without prior dust corrections. We use the dust-attenuated luminosity () and equivalent width () from SPHINX, calculated as described in Sect. 2, and fit the same linear relations between the SFR, , and as in Sect. 3. We present the resulting SFR() calibrations in App. B. The calibration that depends on only (Eq. 4; Fig. 5, cyan) yields dex, while the calibration that additionally depends on (Eq. 5; Fig. 5, magenta) yields dex. In both cases, the RMSE is significantly higher compared to the RMSEs calculated using the calibrations that are based on the intrinsic properties (see Sect. 3). We also find that the SFRs are significantly underestimated at high luminosities (), with Eq. 4 and Eq. 5 producing the median and , respectively. This can be explained by the fact that high- galaxies in SPHINX have, on average, very high attenuation values () which we are unable to fully correct for while simultaneously trying to model the SFRs of low- galaxies. However, we caution that such are not reported in high-redshift observations (see Sect. 2) and thus making accurate SFR estimates based on the dust-attenuated might still be possible in practice.
We also test whether the Balmer decrement (; hereafter ), whose measurements are commonly used to calculate the dust corrections for , can be a useful parameter to include in the SFR() calibration. Interestingly, we find that the -based calibration (Eq. 6; Fig. 5, orange) yields dex, which is dex higher than the RMSE calculated using the -based calibration. This indicates that is not as powerful as in constraining the SFRs of SPHINX galaxies. Including both and in the calibration (Eq. 7; see also Fig. 5, light blue) yields the best results ( dex). Observationally, however, the measurements are hampered by slitlosses and the requirement of a broader wavelength coverage. This implies that the benefits of including in the SFR() calibration are likely limited, with remaining a key addition to the calibration.
4.3 Caveats
When using the SFR() calibrations given in Eq. 2 and Eq. 3, it is important to take proper account of the scope of applicability of these calibrations and the assumptions made to derive them. In particular, the selection criteria described in Sect. 2 imply that the new calibrations can be applied to a flux-limited sample when the following conditions are met:
erg s-1.
Extending the calibrations to erg s-1 is already possible with SPHINX but a larger halo catalog compared to the currently available SPDRv1 catalog is needed to create a complete sample down to very low luminosities. At low redshifts (), the SFHs become less bursty and metallicities increase, rendering the classical calibrations more suitable. Finally, at high redshifts (), it remains to be established whether the new calibrations can be used, as the physical conditions of star formation at such redshifts are poorly understood.
We also stress the fact that the new calibrations assume the BPASSv SPS model with a Kroupa-type IMF (see Sect. 2). Figure 1 shows that the choice of the IMF significantly impacts the predicted SFRs, resulting in a systematic error of up to dex. The choice of the SPS model, in particular whether the model includes binary stars or not, results in a smaller systematic error ( dex), possibly because binary stars have little effect on the LyC luminosity in the first 3 Myrs of a stellar population (e.g., Rosdahl et al., 2018), however such errors are still non-negligible. In the future, these systematic errors can be reduced by testing the SPS models and constraining the IMF at , e.g., by means of deep rest-frame UV spectroscopy (e.g., Steidel et al., 2016; Chisholm et al., 2019).
Finally, we caution that SPHINX does not include black hole formation due to its small volume. This implies that the AGN contribution to the emission is neglected in the new SFR() calibrations. Hence, it is necessary to remove the AGN from the sample before applying these calibrations.
5 Summary
In this work, we study the relationship between emission and the SFR in the SPHINX cosmological simulations at . Using the simulated galaxy data from the SPHINX Public Data Release, Version 1 (Katz et al., 2023), we select a sample of star-forming galaxies that are representative of the -emitter population routinely observed with JWST at . We find that the median SFR— relation in SPHINX exhibits a significant evolution (Fig. 1), with a downturn at low luminosities (–) that is consistent with fainter galaxies having low stellar metallicites () and thus being more efficient producers of ionizing photons than brighter, relatively metal-enriched () galaxies.
Due to the metallicity dependence of the SFR— relation, the classical SFR() calibrations (e.g., Kennicutt, 1998; Eq. 1), on average, overestimate the SFRs of faint galaxies in SPHINX by dex (Fig. 3, black). In Eq. 2, we propose a new SFR() calibration that depends on the intrinsic luminosity only and reduces the error in the predicted SFRs by dex (Fig. 3, cyan). In Eq. 3, we propose another SFR() calibration that additionally depends on the equivalent width, a parameter that is sensitive to the stellar metallicity and age (Fig. 2). This calibration further improves the SFR estimate, reducing the error by a total of dex (Fig. 3, magenta). The new SFR() calibrations affect the galaxy population statistics at high redshifts (): the inferred cosmic star-formation density descreases by %, and the slope of the star-formation main sequence increases by (Fig. 4).
Acknowledgements.
We thank Thibault Garel, Pascal Oesch, Irene Shivaei, Charlotte Simmonds, Andrew Hopkins, Daniel Schaerer, and Rashmi Gottumukkala for useful comments and productive discussions. We gratefully acknowledge support from the CBPsmn (PSMN, Pôle Scientifique de Modélisation Numérique) of the ENS de Lyon for the computing resources. Funded by the European Union (ERC, AGENTS, 101076224). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work made extensive use of several open-source software packages, and we gratefully acknowledge the efforts of their authors: numpy (Harris et al., 2020), astropy (Astropy Collaboration et al., 2022), matplotlib (Hunter, 2007), ipython (Perez & Granger, 2007), and scikit-learn (Pedregosa et al., 2011).References
- Asada et al. (2024) Asada, Y., Sawicki, M., Abraham, R., et al. 2024, MNRAS, 527, 11372
- Astropy Collaboration et al. (2022) Astropy Collaboration, Price-Whelan, A. M., Lim, P. L., et al. 2022, ApJ, 935, 167
- Behroozi et al. (2019) Behroozi, P., Wechsler, R. H., Hearin, A. P., & Conroy, C. 2019, MNRAS, 488, 3143
- Bouché et al. (2010) Bouché, N., Dekel, A., Genzel, R., et al. 2010, ApJ, 718, 1001
- Bouwens et al. (2015) Bouwens, R. J., Illingworth, G. D., Oesch, P. A., et al. 2015, ApJ, 803, 34
- Carvajal-Bohorquez et al. (2025) Carvajal-Bohorquez, C., Ciesla, L., Laporte, N., et al. 2025, arXiv e-prints, arXiv:2507.13160
- Chisholm et al. (2019) Chisholm, J., Rigby, J. R., Bayliss, M., et al. 2019, ApJ, 882, 182
- Chruślińska et al. (2024) Chruślińska, M., Pakmor, R., Matthee, J., & Matsuno, T. 2024, A&A, 686, A186
- Clarke et al. (2024) Clarke, L., Shapley, A. E., Sanders, R. L., et al. 2024, ApJ, 977, 133
- Cole et al. (2025) Cole, J. W., Papovich, C., Finkelstein, S. L., et al. 2025, ApJ, 979, 193
- Covelo-Paz et al. (2025) Covelo-Paz, A., Giovinazzo, E., Oesch, P. A., et al. 2025, A&A, 694, A178
- Cullen et al. (2019) Cullen, F., McLure, R. J., Dunlop, J. S., et al. 2019, MNRAS, 487, 2038
- Curti et al. (2024) Curti, M., Maiolino, R., Curtis-Lake, E., et al. 2024, A&A, 684, A75
- Davies et al. (2019) Davies, L. J. M., Lagos, C. d. P., Katsianis, A., et al. 2019, MNRAS, 483, 1881
- De Looze et al. (2014) De Looze, I., Cormier, D., Lebouteiller, V., et al. 2014, A&A, 568, A62
- Endsley et al. (2024) Endsley, R., Stark, D. P., Whitler, L., et al. 2024, MNRAS, 533, 1111
- Ferland et al. (2017) Ferland, G. J., Chatzikos, M., Guzmán, F., et al. 2017, Rev. Mexicana Astron. Astrofis., 53, 385
- Flores Velázquez et al. (2021) Flores Velázquez, J. A., Gurvich, A. B., Faucher-Giguère, C.-A., et al. 2021, MNRAS, 501, 4812
- Fu et al. (2025) Fu, S., Sun, F., Jiang, L., et al. 2025, arXiv e-prints, arXiv:2503.03829
- Götberg et al. (2017) Götberg, Y., de Mink, S. E., & Groh, J. H. 2017, A&A, 608, A11
- Gunawardhana et al. (2013) Gunawardhana, M. L. P., Hopkins, A. M., Bland-Hawthorn, J., et al. 2013, MNRAS, 433, 2764
- Hao et al. (2011) Hao, C.-N., Kennicutt, R. C., Johnson, B. D., et al. 2011, ApJ, 741, 124
- Harris et al. (2020) Harris, C. R., Millman, K. J., van der Walt, S. J., et al. 2020, Nature, 585, 357
- Hayes et al. (2010) Hayes, M., Schaerer, D., & Östlin, G. 2010, A&A, 509, L5
- Hopkins et al. (2001) Hopkins, A. M., Connolly, A. J., Haarsma, D. B., & Cram, L. E. 2001, AJ, 122, 288
- Hunter (2007) Hunter, J. D. 2007, Computing in Science & Engineering, 9, 90
- Iyer et al. (2020) Iyer, K. G., Tacchella, S., Genel, S., et al. 2020, MNRAS, 498, 430
- Katz et al. (2023) Katz, H., Rosdahl, J., Kimm, T., et al. 2023, The Open Journal of Astrophysics, 6, 44
- Kennicutt & Evans (2012) Kennicutt, R. C. & Evans, N. J. 2012, ARA&A, 50, 531
- Kennicutt (1998) Kennicutt, Jr., R. C. 1998, ARA&A, 36, 189
- Kroupa (2001a) Kroupa, P. 2001a, MNRAS, 322, 231
- Kroupa (2001b) Kroupa, P. 2001b, MNRAS, 322, 231
- Leitherer et al. (1999) Leitherer, C., Schaerer, D., Goldader, J. D., et al. 1999, ApJS, 123, 3
- Lilly et al. (2013) Lilly, S. J., Carollo, C. M., Pipino, A., Renzini, A., & Peng, Y. 2013, ApJ, 772, 119
- Lin et al. (2025) Lin, X., Egami, E., Sun, F., et al. 2025, arXiv e-prints, arXiv:2504.08028
- Ly et al. (2007) Ly, C., Malkan, M. A., Kashikawa, N., et al. 2007, ApJ, 657, 738
- Mascia et al. (2024) Mascia, S., Pentericci, L., Calabrò, A., et al. 2024, A&A, 685, A3
- Matthee et al. (2022) Matthee, J., Feltre, A., Maseda, M., et al. 2022, A&A, 660, A10
- Matthee & Schaye (2019) Matthee, J. & Schaye, J. 2019, MNRAS, 484, 915
- Meyer et al. (2024) Meyer, R. A., Oesch, P. A., Giovinazzo, E., et al. 2024, MNRAS, 535, 1067
- Michel-Dansac et al. (2020) Michel-Dansac, L., Blaizot, J., Garel, T., et al. 2020, A&A, 635, A154
- Nagaraj et al. (2023) Nagaraj, G., Ciardullo, R., Bowman, W. P., Lawson, A., & Gronwall, C. 2023, The Astrophysical Journal, 943, 5
- Naidu et al. (2024) Naidu, R. P., Matthee, J., Kramarenko, I., et al. 2024, arXiv e-prints, arXiv:2410.01874
- Noeske et al. (2007) Noeske, K. G., Weiner, B. J., Faber, S. M., et al. 2007, ApJ, 660, L43
- Pedregosa et al. (2011) Pedregosa, F., Varoquaux, G., Gramfort, A., et al. 2011, Journal of Machine Learning Research, 12, 2825
- Perez & Granger (2007) Perez, F. & Granger, B. E. 2007, Computing in Science and Engineering, 9, 21
- Pirie et al. (2024) Pirie, C. A., Best, P. N., Duncan, K. J., et al. 2024, arXiv e-prints, arXiv:2410.11808
- Rosdahl et al. (2013) Rosdahl, J., Blaizot, J., Aubert, D., Stranex, T., & Teyssier, R. 2013, MNRAS, 436, 2188
- Rosdahl et al. (2022) Rosdahl, J., Blaizot, J., Katz, H., et al. 2022, MNRAS, 515, 2386
- Rosdahl et al. (2018) Rosdahl, J., Katz, H., Blaizot, J., et al. 2018, MNRAS, 479, 994
- Salpeter (1955) Salpeter, E. E. 1955, ApJ, 121, 161
- Scholte et al. (2025) Scholte, D., Cullen, F., Carnall, A. C., et al. 2025, arXiv e-prints, arXiv:2502.10499
- Shim et al. (2009) Shim, H., Colbert, J., Teplitz, H., et al. 2009, ApJ, 696, 785
- Sobral et al. (2013) Sobral, D., Smail, I., Best, P. N., et al. 2013, MNRAS, 428, 1128
- Sparre et al. (2015) Sparre, M., Hayward, C. C., Springel, V., et al. 2015, MNRAS, 447, 3548
- Stanway & Eldridge (2018) Stanway, E. R. & Eldridge, J. J. 2018, MNRAS, 479, 75
- Steidel et al. (2016) Steidel, C. C., Strom, A. L., Pettini, M., et al. 2016, ApJ, 826, 159
- Sun et al. (2025) Sun, F., Wang, F., Yang, J., et al. 2025, ApJ, 980, 12
- Tacchella et al. (2020) Tacchella, S., Forbes, J. C., & Caplar, N. 2020, MNRAS, 497, 698
- Teyssier (2002) Teyssier, R. 2002, A&A, 385, 337
- Theios et al. (2019) Theios, R. L., Steidel, C. C., Strom, A. L., et al. 2019, ApJ, 871, 128
- Wan et al. (2025) Wan, J. T., Tacchella, S., D’Eugenio, F., Johnson, B. D., & van der Wel, A. 2025, arXiv e-prints, arXiv:2504.05281
- Wang & Lilly (2020) Wang, E. & Lilly, S. J. 2020, ApJ, 892, 87
- Weibel et al. (2024) Weibel, A., Oesch, P. A., Barrufet, L., et al. 2024, MNRAS, 533, 1808
- Zavala et al. (2021) Zavala, J. A., Casey, C. M., Manning, S. M., et al. 2021, ApJ, 909, 165
Appendix A Timescales of star formation traced by
It is commonly assumed in the literature that traces star formation over the last Myr. To test this, we use the SFHs published as part of SPDRv1 and calculate the average of the SFR over the last Myr. We then calculate the correlation coefficients between and the SFRs averaged over different timescales. We find that Pearson’s is highest between Myr (), with a maximum at Myr (; Fig. 6, dashed line). Similarly, Kendall’s peaks at Myr (; Fig. 6, solid line). This suggests that a slightly shorter timescale (– Myr) might be more appropriate for . Nevertheless, for consistency with previous works, we use the SFRs averaged over Myr throughout this paper.
Appendix B SFR() calibrations for the dust-attenuated line measurements