### v2

# as a tracer of star formation in the SPHINX cosmological simulations

as a tracer of star formation in the SPHINX cosmological simulations
The emission line in galaxies is a powerful tracer of their recent star formation activity. With the advent of JWST, we are now able to routinely observe in galaxies at high redshift () and thus measure their star formation rates (SFRs). However, using classical SFR() calibrations to derive the SFRs leads to biased results because high-redshift galaxies are commonly characterized by low metallicities and bursty star formation histories, affecting the conversion factor between the luminosity () and the SFR. In this work, we develop a set of new SFR() calibrations that allow us to predict the SFRs of -emitters at with minimal error. We use the SPHINX cosmological simulations to select a sample of star-forming galaxies representative of the -emitter population observed with JWST. We then derive linear corrections to the classical SFR() calibrations, taking into account variations in the physical properties (e.g., stellar metallicities) among individual galaxies. We obtain two new SFR() calibrations that, compared to the classical calibrations, reduce the root mean squared error (RMSE) in the predicted SFRs by dex and dex, respectively. Using the recent JWST NIRCam/grism observations of -emitters at , we show that the new calibrations affect the high-redshift galaxy population statistics: (i) the estimated cosmic star formation rate density decreases by %, and (ii) the observed slope of the star formation main sequence increases by .
Key Words.:
galaxies: fundamental parameters – galaxies: evolution – galaxies: formation – galaxies: high-redshift – galaxies: star formation1 Introduction
One of the most fundamental and basic properties of galaxies is the rate at which new stars form, the star formation rate (SFR). The SFR is sensitive to the interplay between the gas inflow and outflow rates of galaxies, and the efficiency at which this gas is converted into stars (e.g., Bouché et al., 2010; Lilly et al., 2013; Behroozi et al., 2019). The SFR can fluctuate on different timescales (e.g., Sparre et al., 2015; Iyer et al., 2020; Flores Velázquez et al., 2021), which range from dynamical timescales of molecular clouds (s of Myrs), through galaxy merger timescales ( Myr), to long (Gyr-scale) coherences in accretion rates (e.g., Matthee and Schaye, 2019; Wang and Lilly, 2020; Tacchella et al., 2020; Wan et al., 2025). Observationally, SFRs can be measured using various diagnostics (e.g., Kennicutt and Evans, 2012; De Looze et al., 2014; Davies et al., 2019) that are either based on the light from massive, short-lived stars (e.g., the rest-frame UV), or processed radiation (e.g., nebular lines in HII regions or dust emission). Each of these diagnostics probes different timescales of star formation and is affected by dust attenuation differently.
At high redshifts (), the rest-frame UV emission used to be the most sensitive SFR indicator that could be measured for large galaxy samples, tracing star formation on Myr timescales (e.g., Bouwens et al., 2015). However, dust corrections in the rest-frame UV are significant, with potentially up to 50% of the cosmic star formation rate density () at being obscured by dust (e.g., Zavala et al., 2021; Sun et al., 2025). This emphasizes the great importance of alternative SFR indicators that are less sensitive to dust attenuation, such as the emission. The photons are emitted in the rest-frame optical ( Å), where dust corrections are smaller than in the rest-frame UV due to the negative slope of the dust attenuation curve (however, we caution that there is growing evidence of flatter attenuation curves at fixed or stellar mass at high redshifts compared to the local Universe; e.g., Shivaei et al., 2025; Fisher et al., 2025). The luminosity of the line is directly proportional to the luminosity of ionizing photons that are predominantly emitted by very young stellar populations, making sensitive to star formation on shorter timescales ( Myr; e.g., Leitherer et al., 1999). As the JWST can now measure the distribution of luminosities out to based on spectroscopy and narrow-band imaging (e.g., Clarke et al., 2024; Pirie et al., 2025; Covelo-Paz et al., 2025; Fu et al., 2025), it is timely to reassess the role of the H emission as a tracer of the SFR.
The conversion factor between the H luminosity111Throughout this work, we assume a negligible contribution of active galactic nuclei to the H emission. and the SFR primarily depends on assumptions on the properties of stellar populations, in particular, the star formation history (SFH), the initial mass function (IMF), binarity, and stellar metallicity (e.g, Wilkins et al., 2019). The SFH and IMF are relevant because the ionizing photon emissivity from a single stellar population drops sharply after a few Myr when massive, short-lived O stars leave the main sequence. Binary interactions can extend these ionizing lifetimes (Götberg et al., 2017), smoothing the otherwise steep drop-off after a few Myrs. At fixed stellar mass and age, the metallicity affects the effective temperature of stellar atmospheres due to line blanketing effects, affecting the ionizing luminosity as well.
The classical SFR() calibration proposed by Kennicutt (1998) (see also Kennicutt and Evans 2012) assumes a constant SFH over Myr and solar metallicity, based on the typical conditions in galaxies in the low-redshift Universe. However, various spectroscopic observations of galaxies at high redshifts suggest that their SFHs are bursty (e.g., Asada et al., 2024; Endsley et al., 2024; Cole et al., 2025; Carvajal-Bohorquez et al., 2025). Further, given the gas-phase oxygen abundances are typically measured to be % solar (Curti et al., 2024; Scholte et al., 2025), and early galaxies are expected to be -enhanced due to the rapid formation timescales (e.g., Chruślińska et al., 2024), the iron-peak element abundances are likely to be low. This implies hotter stars and a more efficient photon production at a fixed SFR (e.g., Theios et al., 2019) than at low redshift.
The aim of this paper is to revisit the classical SFR() calibrations to predict the SFRs of galaxies at high redshift (). Instead of using idealistic models with constant SFHs and fixed stellar metallicities, we use simulated galaxies and their modeled luminosities from the SPHINX cosmological radiation-hydrodynamical simulation (Rosdahl et al., 2018; Katz et al., 2023). These cosmological simulations self-consistently model SFHs and metal enrichment processes and are calibrated to match various observables at high redshifts (). The simulated volume matches the typical volumes of deep JWST surveys well, which enables a high resolution that can capture SFR variations on short timescales ( Myr) and approximate the distribution of gas in the interstellar medium relevant for radiative transfer effects.
In Sect. 2, we briefly describe the SPHINX simulations and explain how the H luminosities are derived. In Sect. 3, we present a set of simulation-inferred SFR() calibrations that predict the SFRs of SPHINX galaxies with minimal error. Finally, we discuss the implications of our work in Sect. 4 and summarize in Sect. 5.
2 Methods
In this work, we use the simulated galaxy data from SPHINX, a suite of cosmological radiation-hydrodynamic simulations (Rosdahl et al., 2018). SPHINX leverages the adaptive mesh refinement code RAMSES-RT (Rosdahl et al., 2013; Teyssier, 2002) to simulate tens of thousands of galaxies in the epoch of reionization. The ISM is resolved down to co-moving pc, corresponding to a physical resolution of pc at . The dark matter halos are resolved down to the atomic cooling threshold ( M⊙). Hydrodynamics, non-equilibrium thermochemistry, and radiative transfer of Lyman-continuum radiation (LyC; Å) are performed self-consistently on the fly, with subgrid implementations for gas cooling, star formation, metal enrichment, and feedback. The on-the-fly radiative emission and propagation of LyC photons make SPHINX an excellent model for accurately predicting how emission traces star formation in the simulated galaxies.
Our analysis is based on the data from the SPHINX20 Public Data Release, Version 1 (SPDRv1; Katz et al., 2023). SPHINX20 is the largest simulation of the SPHINX suite, with a periodic box with volume of ( cMpc)3 simulated down to a redshift of (Rosdahl et al., 2022). SPHINX20 uses the BPASSv stellar population synthesis (SPS) model (Stanway and Eldridge, 2018), which adopts a Kroupa (2001) IMF with an upper mass cutoff of 100 M⊙ and a slope of from to M⊙ and from to M⊙. The intrinsic Balmer line luminosities are generated by interpolating the CLOUDY v17.03 models (Ferland et al., 2017) for gas cells that host stellar particles with unresolved Stromgren spheres, or by using the non-equilibrium ionizing fractions from the simulation otherwise. For each halo, the total intrinsic line emission is then the sum of the line luminosities of all gas cells within the virial radius of the halo. The equivalent widths are calculated by taking the stellar and nebular continuum emission into account, generated as detailed in Katz et al. (2023).
We post-process the line and continuum radiation, subject to absorption and scattering by dust, using the radiative transfer code RASCAS (Michel-Dansac et al., 2020). Dust is assigned to the gas cells using the Small Magellanic Cloud (SMC) dust model from Laursen et al. (2009). We note that the dust attenuation curve at high redshifts likely has a steeper slope than the SMC slope (e.g., Shivaei et al., 2025; Markov et al., 2025), possibly due to changes in the dust grain properties. Next, 200,000 photon packets are probabilistically distributed to the gas cells, with the initial wavelengths placed at the line centers, thermally broadened, and shifted according to the bulk velocity of the cell. Likewise, 200,000 photon packets are probabilistically distributed to the star particles (gas cells) based on the luminosities of the stellar (nebular) continuum at the line centers. The observed line luminosities and equivalent widths are then obtained by computing the escape fractions of each component (for details, see Katz et al., 2023).
Due to the computation time and data storage limitations, SPDRv1 only includes SPHINX20 galaxies with M⊙ yr-1, where is the SFR averaged over Myr. This excludes objects that are likely too faint to be observed with JWST (Katz et al., 2023), unless lensing is considered (e.g., Bezanson et al., 2024; Naidu et al., 2024). However, an SFR cut also leads to an incompleteness in the SFR— relation at low luminosities, which can bias our results. We therefore limit our selection of galaxies from SPDRv1 to those with an observed luminosity () greater than erg s-1 (here, we use the simulation-based luminosities calculated as described above). This value is motivated by the recent LF measurements from JWST NIRCam/grism, which reach similar luminosities at the faint end (e.g., Fu et al., 2025). Crucially, this selection yields a complete -emitter sample, which we check by running RASCAS on all SPHINX20 halos whose intrinsic luminosity () exceeds the luminosity cut. We also try erg s-1 and erg s-1 cuts, finding little impact on our results. With this additional luminosity cut, our final sample comprises , 66, 48, 36, 22, 17, and 13 galaxies at , 5, 6, 7, 8, 9, and 10, respectively, with a total of galaxies.
The SN feedback model in SPHINX is calibrated to match the UV luminosity function (UVLF) at (see App. C in Rosdahl et al., 2018). A possible caveat is that the stellar mass-halo mass relation in SPHINX is – dex higher than predicted by the Behroozi et al. (2019) model, for example, which matches a broad array of observational data, including stellar mass functions at – and median UV—stellar mass relations at – (see Fig. 4 in Katz et al., 2023). Furthermore, the stellar mass function (SMF) in SPHINX is significantly overestimated compared to the SMF inferred from the JWST/NIRCam observations at – (Weibel et al., 2024). If the stellar masses of SPHINX galaxies are overestimated, the SFRs should also be overestimated on average. This implies that the amount of dust in the simulation should be higher than in the real Universe in order to match the UVLF. The excess of dust would then explain the fact that SPHINX predicts lower and luminosities at fixed stellar mass than those measured by JWST (e.g., Covelo-Paz et al., 2025; Meyer et al., 2024). Although the possibility of SPHINX galaxies having too much dust does not affect our main results as they are based on the intrinsic properties (Sect. 3), it will become important later for understanding the errors in the SFR() calibrations that use the observed (Sect. 4).
3 Results
In this section, we investigate the intrinsic (i.e., without taking dust attenuation into account) SFR— relation in the SPHINX20 cosmological simulation. We compare this relation to the SFR() calibrations from the literature and study the effect of the variations in the physical properties of individual galaxies (metallicity, stellar age, etc.) on the accuracy of the predicted SFRs. Based on this analysis, we develop a set of new SFR() calibrations that minimize the error in -derived SFRs at high redshift.
Unless otherwise specified, we use the SFRs averaged over the last Myr of the SFH. This choice is motivated by the fact that nebular emission lines such as trace stars with lifetimes of –10 Myr (e.g., Kennicutt and Evans, 2012). This is further supported by the simulation itself, where the correlation coefficient between the SFR and peaks between Myr (–; see App. A).
3.1 The SFR— relation in SPHINX
Figure 1 shows the intrinsic SFR— relation for the galaxy sample selected from the SPHINX20 simulation as described in Sect. 2. In the same figure, we show different SFR() calibrations proposed in the literature. These include the original K98 calibration (), the K98 calibration converted to the Kroupa IMF (Kroupa, 2001; ), and two calibrations from Theios et al. (2019, hereafter T19) calculated for a Kroupa-type IMF and solar metallicity (), and the same IMF but subsolar metallicity (; ), respectively. Here, is the conversion factor between and SFR, defined as follows:
Figure 1 reveals that a constant cannot provide an accurate estimate of the SFRs of SPHINX galaxies across the full range. Specifically, the T19 calibration that assumes (the dashed blue line in Fig. 1) matches the median SFR— relation in SPHINX at high luminosities well (). In contrast, the same SFR() calibration at low luminosities () overestimates the SFRs by – dex. Decreasing would result in a better match at low luminosities, but the SFRs at high luminosities would then be underestimated.
We note that the median SFR— relation in SPHINX at low luminosities lies between the two T19 calibrations that differ only in the assumption on metallicity ( vs. ; indicated by the solid and dashed blue lines in Fig. 1, respectively). Motivated by this result, we test whether the variations in metallicity might be a key factor driving the evolution of the median SFR— relation in SPHINX. We split our sample into low- and high- groups ( and , respectively), and calculate the median stellar metallicity222Hereafter, stellar metallicities (and ages) of SPHINX galaxies are weighted by the LyC luminosity. in each group. We obtain and in the low- and high- groups, respectively. In other words, brighter galaxies are more metal-rich on average, which reflects the mass-metallicity relation (MZR) in SPHINX (for a discussion, see Katz et al., 2023), and likely causes an upturn of the median SFR— relation at high luminosities (see also Fig. 1, where the markers are color-coded by metallicity).
Figure 1 further demonstrates that the original K98 calibration (indicated by the dotted orange line) overestimates the SFRs by – dex compared to SPHINX. One of the main reasons for this discrepancy is the difference in the IMFs: K98 uses the Salpeter (1955) IMF, whereas SPHINX uses a Kroupa-type IMF. After rescaling to the Kroupa IMF, the K98 calibration still predicts – dex higher SFRs than the simulation (the dashed orange line in Fig. 1). This difference cannot be eliminated even when using the T19 calibration which, unlike the K98 calibration, assumes the same stellar population synthesis model as SPHINX (i.e., BPASSv; the dashed blue line in Fig. 1). Therefore, the only way to fully reconcile the K98 calibration and the simulation data is to assume a sub-solar metallicity, for some galaxies as low as (the solid blue line in Fig. 1). As discussed in the introduction, a subsolar metallicity like this agrees well with the current estimates of the gas-phase metallicity and also with the direct estimates of stellar metallicities of galaxies at (Steidel et al., 2016; Cullen et al., 2019), although we caution that these were derived assuming constant SFHs, which might impact the results (Matthee et al., 2022).
Figure 1 also reveals that the SFR— relation in SPHINX exhibits significant scatter around the median relation (hereafter, ). In particular, we find that decreases from dex at to dex at , and has a median dex. This scatter is neglected by the existing SFR() calibrations, leading to errors in the predicted SFRs of individual galaxies. To quantify this effect, we introduce the SFR bias factor, defined as the ratio of the SFR predicted from to the true SFR directly tracked by the simulation: . Here, we use the T19 calibration that assumes to calculate SFR() as this calibration provides the best match to the SPHINX data (see Fig. 1). We note that using a different calibration would only change by a constant factor (in logarithmic scale) without affecting any trends observed between and the physical properties of SPHINX galaxies.
We find that decreases with stellar metallicity (, ; Fig. 2, top panel), in agreement with our earlier analysis of the median SFR— relation in SPHINX. However, we also report a significant scatter in at fixed : dex. This indicates that metallicity alone does not cause the variations in the SFRs of SPHINX galaxies with respect to the SFRs predicted from . In particular, we find that shows a negative correlation with stellar age (, ; Fig. 2, bottom panel), with an average scatter of dex. Notably, galaxies with young stellar populations ( Myr) exhibit high (– dex). Even if we use the best-fit SFR() calibration in the form of Eq. 1 (), still reaches dex at – Myr. This bias can be explained by the fact that the existing SFR() calibrations assume constant star formation over a long timescale (typically Myr or Gyr; e.g., Hao et al., 2011). This assumption breaks down in the simulation, where galaxies exhibit very bursty and stochastic SFHs (Katz et al., 2023).
We also examine the relationship between and stellar mass and find a negative correlation (, ), likely driven by the positive scaling of stellar mass with metallicity (see Katz et al., 2023 for the discussion of the MZR in SPHINX). Finally, we check the dependence of on the escape fraction of ionizing radiation () and find a nearly flat trend (, ). This result is expected because the strength of nebular emission lines is inversely proportional to the escape fraction of ionizing photons, , and SPHINX galaxies have low (%) on average. This agrees with high-redshift observations, where is also estimated to be low (%; e.g., Mascia et al., 2024).
3.2 Simulation-informed SFR() calibrations
Using the insights gained from our analysis of the SFR— relation in SPHINX, we discuss possible improvements to the existing SFR() calibrations that take the form of Eq. 1. For example, stellar metallicity could be a useful parameter to include in the calibration to reduce the error in the predicted SFR (see Fig. 2). However, measuring stellar metallicity is difficult in practice, and we therefore only use quantities that can be measured robustly from galaxy spectra. Specifically, we run an ordinary least-squares (OLS) model similar to Eq. 1, but allowing the coefficient to vary. We use the randomly selected of the sample to train the model and the remaining to validate its performance. The resulting SFR() calibration can be written as follows:
where the last term is the effective correction to Eq. 1. The root mean squared error (RMSE) calculated on the training and test datasets is identical ( dex), indicating that the model performs well on unseen data. Compared to the T19 calibration () which yields dex calculated on the full dataset, the RMSE is reduced by dex (see also Fig. 3).
The SFR() calibration given in Eq. 2 corrects the SFRs by a factor that depends on alone, which means that is neglected even when using the updated calibration. To further improve on this, we add a correction to Eq. 2 that depends on the equivalent width (). This choice is motivated by the fact that traces stellar metallicity and age, which both drive (Fig. 2). In addition, is only weakly sensitive to dust attenuation, which implies that dust corrections affect the results less than when other observables are used (e.g., ). We therefore run an OLS model that includes , using the same training sample as in Eq. 2. The resulting SFR() calibration can be written as follows:
where the last two terms are the effective corrections to Eq. 1. The new calibration yields dex on the training and test datasets, which is dex lower than the RMSE calculated using Eq. 2, or dex lower than the RMSE calculated using the T19 calibration (). Figure 3 compares these calibrations as a function of . In particular, this plot shows that the median in the bins is within dex across the entire range of luminosities () when using Eq. 2 or Eq. 3. This is an improvement over the T19 calibration, which overpredicts the SFRs by dex in galaxies with low to medium luminosities (). The standard deviation of in the bins (hereafter ) is nearly identical between the T19 calibration and the SFR() calibration given in Eq. 2 ( dex), but it decreases by dex ( dex) for Eq. 3, especially at medium luminosities ().
We also run a similar model in which we replace with , which is a proxy for the gas-phase (and, to a weaker extent, stellar) metallicity. The resulting calibration shows the same performance as Eq. 3 ( dex). Given the broader wavelength coverage required to measure compared to , we find Eq. 3 to be more practical when applied to observations and hence use it as our best-performance calibration for the remainder of this paper.
4 Discussion
4.1 Implications
The line has long been known as a reliable indicator of the SFR as it traces the ionizing photon luminosity from massive (), short-lived ( Myr) stars (e.g., Kennicutt, 1998). Over the past decades, has been widely employed in numerous spectroscopic (e.g., Shim et al., 2009; Gunawardhana et al., 2013; Nagaraj et al., 2023) and narrow-band imaging (e.g., Ly et al., 2007; Hayes et al., 2010; Sobral et al., 2013) surveys to measure the cosmic SFH up to . With the recent advent of sensitive near-IR spectroscopy at – m, these measurements have been extended to –, in particular thanks to the JWST/NIRCam slitless spectroscopic observations of large samples of -emitters at the same redshifts (e.g., Covelo-Paz et al., 2025; Fu et al., 2025; Lin et al., 2026).
As observations become more accessible at high redshift, it is crucial to revisit the classical SFR() calibrations (e.g., Kennicutt, 1998) that are commonly used at . Based on a realistic galaxy simulation, we have shown that these calibrations bias the predicted SFRs under the physical conditions that are characteristic of high-redshift galaxies observed with JWST (i.e., a low metallicity and bursty star formation), potentially also affecting population statistics. To quantify this effect, we study the effect of the metallicity- and age-dependent SFR bias (see Fig. 2) on the measurements. Specifically, we convert the dust-corrected luminosity function at presented in Schechter form by Fu et al. (2025) into the star formation rate function (SFRF). We perform this conversion by expressing the luminosity via SFR using two SFR— calibrations: the T19 calibration (), and our new SFR() calibration given in Eq. 2, which, by design, takes the dependence of the SFR— relation on metallicity into account. Following previous studies, we then integrate both SFRFs down to yr-1 which corresponds to (Bouwens et al., 2015). The resulting decreases from to , or by %, when using the new calibration. This suggests that the effect of the evolving SFR— relation on is moderate, likely because galaxies in which the SFR bias is highest contribute little to . For example, galaxies with contribute less than % to the total . If we integrate the SFRF only up to , the change in would be more pronounced: would decrease by %.
The SFR bias also affects the correlations between the SFR and other galaxy properties, such as stellar mass. In Fig. 4, we show the SFR— relation, also known as the star formation main sequence (SFMS; Noeske et al., 2007), in the JWST All the Little Things (ALT) survey (Naidu et al., 2024; Di Cesare et al., 2025). We apply different SFR() calibrations to the dust-corrected luminosities and calculate the slope of the SFMS. The slope increases by (%) when we use Eq. 3 instead of the T19 calibration to calculate the SFRs. Equation 2 yields similar results, although they are less significant (). In addition, we find that the scatter in the SFMS decreases by dex (%) when using Eq. 3, most noticeably at the low-mass end ( dex at ). In contrast, Eq. 2 produces almost the same scatter as the T19 calibration ( dex). We note that the SFMS in Fig. 4 is affected by selection effects: some low-mass galaxies are not detected in observations due to their faintness, resulting in a shallow SFMS slope. Temporary cessation or strong suppression of star formation in galaxies during so-called mini-quenching events (e.g., Gelli et al., 2023; Looser et al., 2025) can further amplify this bias. However, a comprehensive discussion of caveats associated with the SFMS slope measurements is beyond the scope of this paper (this discussion can be found in Di Cesare et al., 2025).
4.2 Dust attenuation
In Sect. 3, we ignored the effects of dust attenuation, assuming that the H line measurements (i.e., and ) can be corrected for these effects without errors. This assumption is also usually made for the SFR() calibrations presented in the literature. However, dust attenuation is known to be an important source of systematic error in the -derived SFRs (e.g., Hopkins et al., 2001). This effect likely persists even at as the number of dust in -emitters at these redshifts is found to be non-negligible (e.g., Covelo-Paz et al., 2025 reported the median mag for a sample of -emitters at ).
Here, we test the possibility of deriving SFRs directly from the observed emission without prior dust corrections. We use the dust-attenuated luminosity () and equivalent width () from SPHINX, calculated as described in Sect. 2, and fit the same linear relations between the SFR, , and as in Sect. 3. We present the resulting SFR() calibrations in App. B. The calibration that only depends on (Eq. 4; Fig. 5, cyan) yields dex, while the calibration that additionally depends on (Eq. 5; Fig. 5, magenta) yields dex. In both cases, the RMSE is significantly higher than the RMSEs calculated using the calibrations that are based on the intrinsic properties (see Sect. 3). Moreover, the SFRs are significantly underestimated at high luminosities (), with Eq. 4 and Eq. 5 producing the median and , respectively. This can be explained by the fact that high- galaxies in SPHINX have very high attenuation values () on average, which we are unable to fully correct for while simultaneously trying to model the SFRs of low- galaxies. However, we caution that like this are not reported in high-redshift observations (see Sect. 2), and it might therefore still be possible in practice to make accurate SFR estimates based on the dust-attenuated .
We also test whether the Balmer decrement (; hereafter ), whose measurements are commonly used to calculate the dust corrections for , can be a useful parameter to include in the SFR() calibration. Interestingly, we find that the -based calibration (Eq. 6; Fig. 5, orange) yields dex, which is dex higher than the RMSE calculated using the -based calibration. This indicates that is not as powerful as in constraining the SFRs of the SPHINX galaxies. This result is surprising because is commonly regarded as a robust indicator of dust attenuation (), while is in principle mostly independent of it. However, we still find a negative correlation between and in SPHINX (, ; Fig. 6), possibly because the amount of dust increases with the age of the stellar population, a parameter to which is sensitive. Further, unlike , correlates with stellar metallicity, allowing for an even more accurate prediction of the SFRs.
Inclusion of both and in the calibration (Eq. 7; see also Fig. 5, light blue) provides the best performance among the considered models ( dex). Observationally, however, the measurements are hampered by slit losses and the requirement of a broader wavelength coverage. This implies that the benefits of including in the SFR() calibration are likely limited, with remaining a key addition to the calibration.
4.3 Caveats
When using the SFR() calibrations given in Eq. 2 and Eq. 3, it is important to take proper account of the scope of applicability of these calibrations and the assumptions made to derive them. In particular, the selection criteria described in Sect. 2 imply that the new calibrations can be applied to a flux-limited sample when the following conditions are met:
erg s-1.
It is already possible to extend the calibrations to erg s-1 with SPHINX, but a larger halo catalog than the currently available SPDRv1 catalog is needed to create a complete sample down to very low luminosities. At low redshifts (), the SFHs become less bursty and the metallicities increase, rendering the classical calibrations more suitable. Finally, at high redshift (), it remains to be established whether the new calibrations can be used, as the physical conditions of star formation at these redshifts are poorly understood.
We also stress the fact that the new calibrations assume the BPASSv SPS model with a Kroupa-type IMF (see Sect. 2). Figure 1 shows that the choice of the IMF significantly affects the predicted SFRs, resulting in a systematic error of up to dex (see also Jeřábková et al., 2018, who studied the effect of a non-universal IMF on the SFR— relation). The choice of the SPS model, in particular, whether the model includes binary stars, results in a smaller error ( dex), possibly because binary stars have little effect on the LyC luminosity in the first Myrs of a stellar population when the bulk of ionizing photons is produced (e.g., Rosdahl et al., 2018), but this effect is still non-negligible. In the future, these systematic errors can be reduced by testing the SPS models and constraining the IMF at , for instance, by means of deep rest-frame UV spectroscopy (e.g., Steidel et al., 2016; Chisholm et al., 2019).
5 Summary
In this work, we study the relation between emission and the SFR in the SPHINX cosmological simulations at . Using the simulated galaxy data from the SPHINX Public Data Release, Version 1 (Katz et al., 2023), we select a sample of star-forming galaxies that are representative of the -emitter population routinely observed with JWST at . The median SFR— relation in SPHINX exhibits a downturn at low luminosities (Fig. 1; –). This downturn is consistent with fainter galaxies having lower stellar metallicities () and thus being more efficient producers of ionizing photons than brighter, relatively metal-enriched () galaxies. Furthermore, the SFR— relation exhibits a significant scatter at fixed luminosity, ranging from dex at to dex at (i.e., the scatter is larger at lower luminosities). We demonstrate that this scatter is primarily driven by the variations in the stellar metallicities and ages (Fig. 2), whereas other properties (e.g., and ) play a less significant role.
Because of the metallicity dependence of the SFR— relation, the classical SFR() calibrations (e.g., Kennicutt, 1998; Eq. 1) overestimate the SFRs of faint galaxies in SPHINX on average by dex (Fig. 3, black). In Eq. 2, we propose a new SFR() calibration that depends on the intrinsic luminosity alone and reduces the error in the predicted SFRs by dex (Fig. 3, cyan). To address the scatter in the SFR— relation, we propose another SFR() calibration in Eq. 3. This calibration additionally depends on the equivalent width, a parameter that is sensitive to the stellar metallicity and age (Fig. 2), and further improves the SFR estimate, reducing the error by a total of dex (Fig. 3, magenta). The new SFR() calibrations affect the galaxy population statistics at high redshift (): the inferred decreases by %, and the slope of the SFMS increases by (Fig. 4).
Acknowledgements.
We thank the anonymous referee for the insightful comments that helped improve the manuscript. We also thank Thibault Garel, Pascal Oesch, Irene Shivaei, Charlotte Simmonds, Andrew Hopkins, Daniel Schaerer, and Rashmi Gottumukkala for useful comments and productive discussions. We gratefully acknowledge support from the CBPsmn (PSMN, Pôle Scientifique de Modélisation Numérique) of the ENS de Lyon for the computing resources. Funded by the European Union (ERC, AGENTS, 101076224). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work made extensive use of several open-source software packages, and we gratefully acknowledge the efforts of their authors: numpy (Harris et al., 2020), astropy (Astropy Collaboration et al., 2022), matplotlib (Hunter, 2007), ipython (Perez and Granger, 2007), and scikit-learn (Pedregosa et al., 2011).References
- Bursty star formation and galaxy-galaxy interactions in low-mass galaxies 1 Gyr after the Big Bang. 527 (4), pp. 11372–11392. External Links: Document, 2310.02314, ADS entry Cited by: §1.
- The Astropy Project: Sustaining and Growing a Community-oriented Open-source Project and the Latest Major Release (v5.0) of the Core Package. 935 (2), pp. 167. External Links: Document, 2206.14220, ADS entry Cited by: §5.
- UNIVERSEMACHINE: The correlation between galaxy growth and dark matter halo assembly from z = 0-10. 488 (3), pp. 3143–3194. External Links: Document, 1806.07893, ADS entry Cited by: §1, §2.
- The JWST UNCOVER Treasury Survey: Ultradeep NIRSpec and NIRCam Observations before the Epoch of Reionization. ApJ 974 (1), pp. 92. External Links: Document, 2212.04026, ADS entry Cited by: §2.
- The Impact of Cold Gas Accretion Above a Mass Floor on Galaxy Scaling Relations. 718 (2), pp. 1001–1018. External Links: Document, 0912.1858, ADS entry Cited by: §1.
- UV Luminosity Functions at Redshifts z 4 to z 10: 10,000 Galaxies from HST Legacy Fields. 803 (1), pp. 34. External Links: Document, 1403.4295, ADS entry Cited by: §1, §4.1.
- Stochastic star formation activity of galaxies within the first billion years probed by JWST. A&A 704, pp. A290. External Links: Document, 2507.13160, ADS entry Cited by: §1.
- Constraining the Metallicities, Ages, Star Formation Histories, and Ionizing Continua of Extragalactic Massive Star Populations. ApJ 882 (2), pp. 182. External Links: Document, 1905.04314, ADS entry Cited by: §4.3.
- Trading oxygen for iron. I. The [O/Fe]-specific star formation rate relation of galaxies. 686, pp. A186. External Links: Document, 2308.00023, ADS entry Cited by: §1.
- The Star-forming Main Sequence in JADES and CEERS at z ¿ 1.4: Investigating the Burstiness of Star Formation. 977 (1), pp. 133. External Links: Document, 2406.05178, ADS entry Cited by: §1.
- CEERS: Increasing Scatter along the Star-forming Main Sequence Indicates Early Galaxies Form in Bursts. 979 (2), pp. 193. External Links: Document, 2312.10152, ADS entry Cited by: §1.
- An H view of galaxy buildup in the first 2 Gyr: Luminosity functions at z 4‑6.5 from NIRCam/grism spectroscopy. 694, pp. A178. External Links: Document, 2409.17241, ADS entry Cited by: §1, §2, §4.1, §4.2.
- The VANDELS survey: the stellar metallicities of star-forming galaxies at 2.5 ¡ z ¡ 5.0. 487 (2), pp. 2038–2060. External Links: Document, 1903.11081, ADS entry Cited by: §3.1.
- JADES: Insights into the low-mass end of the mass-metallicity-SFR relation at 3 ¡ z ¡ 10 from deep JWST/NIRSpec spectroscopy. 684, pp. A75. External Links: Document, 2304.08516, ADS entry Cited by: §1.
- Galaxy And Mass Assembly (GAMA): The sSFR-M∗ relation part I - sSFR-M∗ as a function of sample, SFR indicator, and environment. 483 (2), pp. 1881–1900. External Links: Document, 1811.03712, ADS entry Cited by: §1.
- The applicability of far-infrared fine-structure lines as star formation rate tracers over wide ranges of metallicities and galaxy types. A&A 568, pp. A62. External Links: Document, 1402.4075, ADS entry Cited by: §1.
- The slope and scatter of the star forming main sequence at z~5 : reconciling observations with simulations. arXiv e-prints, pp. arXiv:2510.19044. External Links: Document, 2510.19044, ADS entry Cited by: Figure 4, §4.1.
- The star-forming and ionizing properties of dwarf z 6-9 galaxies in JADES: insights on bursty star formation and ionized bubble growth. 533 (1), pp. 1111–1142. External Links: Document, 2306.05295, ADS entry Cited by: §1.
- The 2017 Release Cloudy. Rev. Mexicana Astron. Astrofis. 53, pp. 385–438. External Links: Document, 1705.10877, ADS entry Cited by: §2.
- REBELS-IFU: dust attenuation curves of 12 massive galaxies at z ≃ 7. MNRAS 539 (1), pp. 109–126. External Links: Document, 2501.10541, ADS entry Cited by: §1.
- The time-scales probed by star formation rate indicators for realistic, bursty star formation histories from the FIRE simulations. 501 (4), pp. 4812–4824. External Links: Document, 2008.08582, ADS entry Cited by: §1.
- Medium-band Astrophysics with the Grism of NIRCam In Frontier Fields (MAGNIF): Spectroscopic Census of H Luminosity Functions and Cosmic Star Formation at z 4.5 and 6.3. 987 (2), pp. 186. External Links: Document, 2503.03829, ADS entry Cited by: §1, §2, §4.1, §4.1.
- Quiescent Low-mass Galaxies Observed by JWST in the Epoch of Reionization. ApJ 954 (1), pp. L11. External Links: Document, 2303.13574, ADS entry Cited by: §4.1.
- Ionizing spectra of stars that lose their envelope through interaction with a binary companion: role of metallicity. 608, pp. A11. External Links: Document, 1701.07439, ADS entry Cited by: §1.
- Galaxy And Mass Assembly: evolution of the H luminosity function and star formation rate density up to z ¡ 0.35. MNRAS 433 (4), pp. 2764–2789. External Links: Document, 1305.5308, ADS entry Cited by: §4.1.
- Dust-corrected Star Formation Rates of Galaxies. II. Combinations of Ultraviolet and Infrared Tracers. 741 (2), pp. 124. External Links: Document, 1108.2837, ADS entry Cited by: §3.1.
- Array programming with NumPy. 585 (7825), pp. 357–362. External Links: Document, Link Cited by: §5.
- The H-alpha luminosity function at redshift 2.2 . A new determination using VLT/HAWK-I. 509, pp. L5. External Links: Document, 0912.3267, ADS entry Cited by: §4.1.
- Toward a Resolution of the Discrepancy between Different Estimators of Star Formation Rate. AJ 122 (1), pp. 288–296. External Links: Document, astro-ph/0103253, ADS entry Cited by: §4.2.
- Matplotlib: a 2d graphics environment. 9 (3), pp. 90–95. External Links: Document Cited by: §5.
- The diversity and variability of star formation histories in models of galaxy evolution. 498 (1), pp. 430–463. External Links: Document, 2007.07916, ADS entry Cited by: §1.
- Impact of metallicity and star formation rate on the time-dependent, galaxy-wide stellar initial mass function. A&A 620, pp. A39. External Links: Document, 1809.04603, ADS entry Cited by: §4.3.
- The SPHINX Public Data Release: Forward Modelling High-Redshift JWST Observations with Cosmological Radiation Hydrodynamics Simulations. 6, pp. 44. External Links: Document, 2309.03269, ADS entry Cited by: §1, §2, §2, §2, §2, §3.1, §3.1, §3.1, §5.
- Star Formation in the Milky Way and Nearby Galaxies. 50, pp. 531–608. External Links: Document, 1204.3552, ADS entry Cited by: §1, §1, §3.
- Star Formation in Galaxies Along the Hubble Sequence. 36, pp. 189–232. External Links: Document, astro-ph/9807187, ADS entry Cited by: §1, §4.1, §4.1, §5.
- On the variation of the initial mass function. 322 (2), pp. 231–246. External Links: Document, astro-ph/0009005, ADS entry Cited by: §2, §3.1.
- Ly Radiative Transfer with Dust: Escape Fractions from Simulated High-Redshift Galaxies. ApJ 704 (2), pp. 1640–1656. External Links: Document, 0907.2698, ADS entry Cited by: §2.
- Starburst99: Synthesis Models for Galaxies with Active Star Formation. 123 (1), pp. 3–40. External Links: Document, astro-ph/9902334, ADS entry Cited by: §1.
- Gas Regulation of Galaxies: The Evolution of the Cosmic Specific Star Formation Rate, the Metallicity-Mass-Star-formation Rate Relation, and the Stellar Content of Halos. 772 (2), pp. 119. External Links: Document, 1303.5059, ADS entry Cited by: §1.
- The Luminosity Function and Clustering of H Emitting Galaxies at z ≍ 46 from a Complete NIRCam Grism Redshift Survey. 997 (2), pp. 207. External Links: Document, 2504.08028, ADS entry Cited by: §4.1.
- JADES: Differing assembly histories of galaxies: Observational evidence for bursty star formation histories and (mini-)quenching in the first billion years of the Universe. A&A 697, pp. A88. External Links: Document, 2306.02470, ADS entry Cited by: §4.1.
- The Luminosity Function and Star Formation Rate between Redshifts of 0.07 and 1.47 for Narrowband Emitters in the Subaru Deep Field. ApJ 657 (2), pp. 738–759. External Links: Document, astro-ph/0610846, ADS entry Cited by: §4.1.
- The evolution of dust attenuation in z 2-12 galaxies observed by JWST. Nature Astronomy 9, pp. 458–468. External Links: Document, 2402.05996, ADS entry Cited by: §2.
- New insight on the nature of cosmic reionizers from the CEERS survey. 685, pp. A3. External Links: Document, 2309.02219, ADS entry Cited by: §3.1.
- Deciphering stellar metallicities in the early Universe: case study of a young galaxy at z = 4.77 in the MUSE eXtremely Deep Field. 660, pp. A10. External Links: Document, 2111.14855, ADS entry Cited by: §3.1.
- The origin of scatter in the star formation rate-stellar mass relation. 484 (1), pp. 915–932. External Links: Document, 1805.05956, ADS entry Cited by: §1.
- JWST FRESCO: a comprehensive census of H + [O III] emitters at 6.8 ¡ z ¡ 9.0 in the GOODS fields. MNRAS 535 (1), pp. 1067–1094. External Links: Document, 2405.05111, ADS entry Cited by: §2.
- RASCAS: RAdiation SCattering in Astrophysical Simulations. A&A 635, pp. A154. External Links: Document, 2001.11252, ADS entry Cited by: §2.
- The hα and [o iii] λ5007 luminosity functions of 1.2 lt; z lt; 1.9 emission-line galaxies from hubble space telescope (hst) grism spectroscopy. The Astrophysical JournalA&AAJApJApJMNRASApJA&AApJApJApJApJMNRASApJApJMNRASarXiv e-printsA&AMNRASMNRASA&AMNRASApJMNRASA&AMNRASPASPA&AarXiv e-printsA&AMNRASApJA&AApJSApJApJMNRASMNRASMNRASApJMNRASMNRASMNRASMNRASApJMNRASApJApJARA&AApJARA&AApJMNRASThe Open Journal of AstrophysicsComputing in Science and EngineeringComputing in Science & EngineeringNatureApJJournal of Machine Learning Research 943 (1), pp. 5. External Links: Document, Link Cited by: §4.1.
- All the Little Things in Abell 2744: 1000 Gravitationally Lensed Dwarf Galaxies at from JWST NIRCam Grism Spectroscopy. pp. arXiv:2410.01874. External Links: Document, 2410.01874, ADS entry Cited by: §2, Figure 4, §4.1.
- Star Formation in AEGIS Field Galaxies since z=1.1: The Dominance of Gradually Declining Star Formation, and the Main Sequence of Star-forming Galaxies. 660 (1), pp. L43–L46. External Links: Document, astro-ph/0701924, ADS entry Cited by: §4.1.
- Scikit-learn: machine learning in Python. 12, pp. 2825–2830. Cited by: §5.
- IPython: A System for Interactive Scientific Computing. 9 (3), pp. 21–29. External Links: Document, ADS entry Cited by: §5.
- The JWST Emission Line Survey (JELS): an untargeted search for H emission line galaxies at z ¿ 6 and their physical properties. 541 (2), pp. 1348–1376. External Links: Document, 2410.11808, ADS entry Cited by: §1.
- RAMSES-RT: radiation hydrodynamics in the cosmological context. MNRAS 436 (3), pp. 2188–2231. External Links: Document, 1304.7126, ADS entry Cited by: §2.
- LyC escape from SPHINX galaxies in the Epoch of Reionization. MNRAS 515 (2), pp. 2386–2414. External Links: Document, 2207.03232, ADS entry Cited by: §2.
- The SPHINX cosmological simulations of the first billion years: the impact of binary stars on reionization. 479 (1), pp. 994–1016. External Links: Document, 1801.07259, ADS entry Cited by: §1, §2, §2, §4.3.
- The Luminosity Function and Stellar Evolution.. 121, pp. 161. External Links: Document, ADS entry Cited by: §3.1.
- The JWST EXCELS survey: probing strong-line diagnostics and the chemical evolution of galaxies over cosmic time using Te-metallicities. 540 (2), pp. 1800–1826. External Links: Document, 2502.10499, ADS entry Cited by: §1.
- Global Star Formation Rate Density over 0.7 ¡ z ¡ 1.9. ApJ 696 (1), pp. 785–796. External Links: Document, 0902.0736, ADS entry Cited by: §4.1.
- The Diversity and Evolution of Dust Attenuation Curves from Redshift z ~1 to 9. arXiv e-prints, pp. arXiv:2509.01795. External Links: Document, 2509.01795, ADS entry Cited by: §1, §2.
- A large H survey at z = 2.23, 1.47, 0.84 and 0.40: the 11 Gyr evolution of star-forming galaxies from HiZELS★. 428 (2), pp. 1128–1146. External Links: Document, 1202.3436, ADS entry Cited by: §4.1.
- The star formation main sequence and stellar mass assembly of galaxies in the Illustris simulation. 447 (4), pp. 3548–3563. External Links: Document, 1409.0009, ADS entry Cited by: §1.
- Re-evaluating old stellar populations. 479 (1), pp. 75–93. External Links: Document, 1805.08784, ADS entry Cited by: §2.
- Reconciling the Stellar and Nebular Spectra of High-redshift Galaxies. 826 (2), pp. 159. External Links: Document, 1605.07186, ADS entry Cited by: §3.1, §4.3.
- A SPectroscopic Survey of Biased Halos in the Reionization Era (ASPIRE): Spectroscopically Complete Census of Obscured Cosmic Star Formation Rate Density at z = 4–6. 980 (1), pp. 12. External Links: Document, 2412.06894, ADS entry Cited by: §1.
- Stochastic modelling of star-formation histories II: star-formation variability from molecular clouds and gas inflow. 497 (1), pp. 698–725. External Links: Document, 2006.09382, ADS entry Cited by: §1.
- Cosmological hydrodynamics with adaptive mesh refinement. A new high resolution code called RAMSES. A&A 385, pp. 337–364. External Links: Document, astro-ph/0111367, ADS entry Cited by: §2.
- Dust Attenuation, Star Formation, and Metallicity in z 2-3 Galaxies from KBSS-MOSFIRE. 871 (1), pp. 128. External Links: Document, 1805.00016, ADS entry Cited by: §1, §3.1.
- Decoding the variability in the star formation histories of z 0.8 galaxies. 539 (4), pp. 2891–2909. External Links: Document, 2504.05281, ADS entry Cited by: §1.
- The Variability of the Star Formation Rate in Galaxies. I. Star Formation Histories Traced by EW(H) and EW(HA). 892 (2), pp. 87. External Links: Document, 1912.06523, ADS entry Cited by: §1.
- Galaxy build-up in the first 1.5 Gyr of cosmic history: insights from the stellar mass function at z 4-9 from JWST NIRCam observations. MNRAS 533 (2), pp. 1808–1838. External Links: Document, 2403.08872, ADS entry Cited by: §2.
- Recalibrating the cosmic star formation history. MNRAS 490 (4), pp. 5359–5365. External Links: Document, 1910.05220, ADS entry Cited by: §1.
- The Evolution of the IR Luminosity Function and Dust-obscured Star Formation over the Past 13 Billion Years. 909 (2), pp. 165. External Links: Document, 2101.04734, ADS entry Cited by: §1.
Appendix A Timescales of star formation traced by
It is commonly assumed in the literature that traces star formation over the last Myr. To test this, we use the SFHs published as part of SPDRv1 and calculate the average of the SFR over the last Myr. We then calculate the correlation coefficients between and the SFRs averaged over different timescales. We find that Pearson’s is highest between Myr (), with a maximum at Myr (; Fig. 7, dashed line). Similarly, Kendall’s peaks at Myr (; Fig. 7, solid line). This suggests that a slightly shorter timescale (– Myr) might be more appropriate for . Nevertheless, for consistency with previous works, we use the SFRs averaged over Myr throughout this paper.
Appendix B SFR() calibrations for the dust-attenuated line measurements